FWD: The number of passengers p and the number of suitcases s that an airplane can carry are related by the equation 180p+60s=3000.If 10 People board the aircraft,how many suitcases can the airplane carry?
Answer by sachi(548)

180p+60s=3000
when p=10
then 180*10+60s=3000
or 60s=3000-1800=1200
or s=1200/60=20
ans